OpenStudy (anonymous):

write the following equation in slope intercept form or point-slope form. the line through the points (2,4) and (4,-5)

OpenStudy (anonymous):

sure. first find the slope using \[m=\frac{y_2-y_1}{x_2-x_1}\] you get \[\frac{-5-4}{4-2}=\frac{-9}{2}\]

OpenStudy (anonymous):

then use the point- slope formula \[y-y_1=m(x-x_1)\] with \[m=-\frac{9}{2},x_1=2,y_1=4\] to get \[y-4=-\frac{9}{2}(x-2)\] \[y-4=-\frac{9}{2}x+9\] \[y=-\frac{9}{2}x+13\]